# Kolmogorov-Arnold Energy Models: Fast and Interpretable Generative Modeling

### v8

@@ Abstract
Learning an energy-based model (EBM) in the latent space of a top-down generative model offers a powerful framework for generation across many data modalities. However, it remains unclear how its interpretability can be used to guide model design, improve generative quality, and reduce training time. Moreover, the reliance on Langevin Monte Carlo (LMC) sampling presents challenges in efficiency and sampling multimodal latent distributions. We propose a novel adaptation of the Kolmogorov-Arnold representation theorem for generative modeling and introduce the Kolmogorov-Arnold Energy Model (KAEM) to take advantage of structural and inductive biases. By constraining the prior to univariate relationships, KAEM enables fast and exact inference via the inverse transform method. With the low dimensionality of the latent space and suitable inductive biases encoded, we demonstrate that importance sampling (IS) becomes a viable, unbiased, and highly efficient posterior sampler. For domains where IS fails, we introduce a strategy based on population-based LMC, decomposing the posterior into a sequence of annealed distributions to improve LMC mixing. KAEM balances common generative modeling trade-offs, offering fast inference, interpretability, and stable training, while being naturally suited to Zettascale Computing hardware.
@@ 1 Introduction
Deep generative modeling is a core subfield of machine learning that enables systems to learn complex data distributions, generate novel data, and support a wide range of tasks including simulation, compression, and representation learning. Two prominent and widely studied approaches are Variational Autoencoders (VAEs) (kingma2022autoencodingvariationalbayes) and Generative Adversarial Networks (GANs) (goodfellow2014generativeadversarialnetworks). Both frameworks assume that each data point is generated from a low-dimensional latent vector via a top-down neural network, where the latent vector is typically drawn from a simple, non-informative prior, such as a uniform or isotropic Gaussian distribution.
Latent energy-based prior models (EBMs) (pang2020learninglatentspaceenergybased) extend this framework by learning an expressive, data-dependent prior directly in the latent space. This prior can be trained jointly with the top-down generator using maximum likelihood estimation (MLE), where the EBM learns by the statistical discrepancy between samples from the prior and posterior distributions, following a paradigm similar to Empirical Bayes. Each learning iteration requires Markov Chain Monte Carlo (MCMC) sampling from both the prior and posterior, specifically the Unadjusted Langevin Algorithm (ULA) (Brooks2011; Rossky1978; Roberts2002), which often offers greater modeling flexibility compared to the amortized inference methods of VAEs.
Despite being extendable to several high-impact research directions, such as multimodal generative modeling (yuan2024learningmultimodallatentgenerative), and interpretable classification, text-to-text, and meta-learning (pang2021latentspaceenergybasedmodeltext), (which involves rapid adaptability and few-shot generalization to out-of-distribution tasks), latent EBMs face critical limitations due to their reliance on MCMC sampling.
Firstly, MCMC is inherently iterative, and algorithms such as ULA require repeated gradient passes through the model, which introduces significant computational overhead during both training and inference. Secondly, ULA is best suited for sampling from smooth, unimodal distributions and can struggle to explore the multimodal latent distributions that are typical in latent variable generative modeling. Moreover, the framework presented by pang2020learninglatentspaceenergybased requires separate ULA samplers for the prior and posterior. This introduces additional experimental complexity since ULA is sensitive to hyperparameter tuning: step sizes that are too small lead to poor mixing and convergence, while larger step sizes can accumulate discretization bias due to its use of the Euler-Maruyama method (Wilkie_2004) to evolve the underlying Langevin equation.
Additionally, it is unclear how the unique separation between the prior and generator networks can be leveraged for interpretability or training efficiency. Unlike the framework introduced by Karniadakis2021, which describes how scientific inductive and learning biases can be embedded into neural networks to reduce their reliance on data and improve training efficiency, generalization, and explainability, (all of which are critical for ensuring trust in Artificial Intelligence), it remains unclear how the EBM prior can be effectively aligned with the posterior to achieve similar effects. Although ideas from mechanistic interpretability, as reviewed by bereska2024mechanisticinterpretabilityaisafety, could potentially be applied, there is currently no established method to leverage the unique strengths of the EBM prior framework for this purpose.
1.1 Inductive biases and fast sampling
To address these limitations, we reimagine the EBM prior into a novel univariate or element-wise model, drawing inspiration from Kolmogorov-Arnold Networks (KANs) (liu2024kankolmogorovarnoldnetworks) and deep spline networks 9264754, which replace the sigmoidal activations used in multilayer perceptrons MURTAGH1991183 with spline interpolations. By representing the latent space with univariate functions and distributions, we enable fast, (single forward pass), and exact prior sampling with the inverse transform method (devroye1986nonuniform) and improve interpretability for domain experts who may not be deeply familiar with machine learning. This allows the EBM prior to be aligned with the posterior before training begins, and for latent distributions to be plotted, viewed, and potentially understood with the symbolic regression approach proposed by liu2024kankolmogorovarnoldnetworks.
@@ 1.1 Inductive biases and fast sampling
In Sec. 3.1, we show that simple datasets can allow for efficient training with importance sampling (IS) (goertzel1949quota). This is a technique for estimating posterior expectations that is otherwise avoided in the literature due to its high variance under severe prior-posterior mismatch and the curse of dimensionality, (dupuis2004importance), which can be avoided in our model due to the low dimensionality of the latent space. Moreover, we show that the trained priors are easy to visualize and recover, offering new opportunities for latent discovery and knowledge transfer to models with different latent structures.
To derive our new model, which we call the Kolmogorov-Arnold Energy Model (KAEM), we first introduce a novel adaptation of the Kolmogorov-Arnold Representation Theorem (KART) (Arnold2009) by reinterpreting its inner functions as Markov kernels between probability spaces. Given its finite and definitive nature, KART can provide structural bias and reduce the reliance of architecture design on intuition and automated tuning.
This reinterpretation allows us to frame inverse transform sampling (ITS) as a measurable transformation applied to uniform noise, thus preserving the deterministic structure of KART while admitting a probabilistic interpretation that aligns with MLE. To address the potential inflexibility of strict adherence to KART and the limitations of using independent energy-based priors, we introduce a scalable strategy based on mixture distributions. Beyond introducing dependencies between latent dimensions to capture inter-dimensional relationships, this also improves sampling efficiency by enabling component-wise sampling.
1.2 Population-based Langevin sampling and Thermodynamic training
However, while ITS enables fast inference, posterior sampling must still be conducted using ULA when the IS estimator is rendered ineffective due to variance. Prior work addressing poor MCMC convergence and multimodal exploration in EBMs has typically relied on diffusion models, such as the hierarchical EBM by cui2024learninglatentspacehierarchical and the diffusion-assisted EBM by zhang2023persistentlytraineddiffusionassistedenergybased.
To maintain prior interpretability and fast inference post-training, we instead introduce posterior annealing as a novel and alternative strategy. This involves decomposing the posterior into a sequence of power posteriors, presented by 2d186535-310b-3428-8c6f-e4ef298483d7, and recasting ULA with population-based MCMC methods (PT; Marinari1992; Hukushima1996).
The training criterion is derived from the discretized Thermodynamic Integral and the Steppingstone Estimator (SE), as presented by CALDERHEAD20094028; Annis2019 to ensure the impact of annealing on learning dynamics and parameter updates. Although this might be less expressive than hierarchical EBMs, it ensures that inductive biases remain easy to embed into the latent prior and interpret after training; and unlike denoising, which scales sequentially, annealing can scale by adding more annealed distributions in parallel.
While ULA is computationally efficient and enables more flexible chain dynamics, it does not preserve detailed balance or reversibility. This can be recovered by incorporating Metropolis-Hastings (MH) adjustments (13ab5b5e-0237-33fb-a7a8-6f6e4e0d4e0f) via the Metropolis-adjusted Langevin Algorithm (MALA) (10.1111/j.1467-9868.2010.00765.x). However, ULA remains the preferred method in our work, motivated by its efficiency and effectiveness, demonstrated by pang2020learninglatentspaceenergybased and further supported by the findings of nijkamp2019learningnonconvergentnonpersistentshortrun; nijkamp2019anatomymcmcbasedmaximumlikelihood, with regards to how short-run, biased MCMC methods can still be highly effective for sampling from EBMs.
MLE and SE are compared in Sec. 3.2. Despite our optimizations, the element-wise and localized nature of KANs complicate their parallelization on GPUs, which are best suited to matrix arithmetic. As a result, forward passes through KAEM and the gradient evaluations required for ULA were found to be slow, necessitating reduced experimental complexity.
1.3 Scope of experiments
@@
KAEM was compiled and differentiated using Julia’s Reactant (Reactantjl2025) and Enzyme (moses2020enzyme; moses2021enzyme_gpu; moses2022enzyme_parallel) packages, which optimizes and lowers Julia code to Multi-Level Intermediate Representation (MLIR) (lattner2020mlircompilerinfrastructureend), thus offering best-in-class performance. However, Gaussian Radial Basis Functions (RBF KANs) (li2024kolmogorovarnoldnetworksradialbasis), were adopted instead of the cubic B-spline bases proposed by liu2024kankolmogorovarnoldnetworks, as RBFs are Reactant-compatible and more efficient for GPU implementation. They have also been proven to provide an optimal solution to the regularized approximation/interpolation problem by 58326.
The experiments in Sec. 3.2 demonstrate that our mixture prior model is able to generate coherent, (although low-quality), samples after a few training epochs on RGB images, while still supporting efficient, exact sampling via component-wise ITS. In contrast, annealing incurred substantial computational overhead, but produced small improvements in sample quality, rendering its recommendation context-dependent. However, this may change as a larger number of power posteriors, longer training durations, and more advanced posterior sampling algorithms become practical with the hardware advances proposed by exa2024. Their architecture enables fine-grained control over data flow and parallelized element-wise nonlinearity which is naturally well suited to supporting KAEM at scale. Their XPUs also provide function/layer fusion capabilities that may improve the potential of univariate priors and adherence to KART.
@@
We present KAEM as an initial step to inspire further adaptations of KART in machine learning. This work aims to lay the groundwork for more exploration amongst the research community, with the far-reaching hope of eventually advancing towards a broader conceptualization: “The Kolmogorov-Arnold Representation Theorem Is All You Need.“
2 The Kolmogorov-Arnold Energy Model
This section introduces the model and its associated sampling procedures. We begin by presenting a novel interpretation of the Kolmogorov-Arnold Representation Theorem (KART) to define the model’s baseline architecture. Our key practical contribution is the introduction of a univariate energy-based prior, detailed in Sec. 2.1. The top-down generator can be instantiated in various forms, including Kolmogorov-Arnold Networks (KANs), which enables strict adherence to KART.
We then propose scalable extensions to the model, including mixture distributions in Sec. 2.3.2, which introduce dependencies between latent dimensions. This represents a novel contribution to the literature on EBM priors. The final sections address prior and posterior sampling: we introduce inverse transform sampling (ITS) as a new approach for efficient EBM prior sampling, and introduce the posterior annealing strategy in Sec. 2.5 to improve multimodal sampling during training.
@@ 2 The Kolmogorov-Arnold Energy Model
The Kolmogorov-Arnold Representation theorem (KART) (Arnold2009) forms the basis of KAEM. A structured proof of the theorem has been provided by dzhenzher2022structuredproofkolmogorovssuperposition.
The Kolmgorov-Arnold Representation theorem 1
@@
Eq. 1 can provide structural bias and guide the design of KAEM’s architecture. To ensure adherence to it, any transformations must be derived to be continuous, measurable, and deterministic. For readers less familiar with the technical background, the two most practically significant equations of Sec. 2.1 are Eqs. 46.
2.1 Univariate energy-based prior

### v9

# Kolmogorov-Arnold Energy Models: Fast, Interpretable Generative Modeling

@@ Abstract
Generative models typically rely on either simple latent priors (e.g., Variational Autoencoders, VAEs), which are efficient but limited, or highly expressive iterative samplers (e.g., Diffusion and Energy-based Models), which are costly and opaque. We introduce the Kolmogorov-Arnold Energy Model (KAEM) to bridge this trade-off and provide a new avenue for latent-space interpretability. Based on a novel interpretation of the Kolmogorov-Arnold Representation Theorem, which provides new structural biases for generative modeling, KAEM imposes a univariate latent structure that enables fast and exact inference via the inverse transform method. With a low-dimensional latent space and appropriate inductive biases, we show that importance sampling becomes a viable, unbiased, and highly efficient posterior inference method. For settings where importance sampling fails, we propose a population-based strategy that decomposes the posterior into a sequence of annealed distributions to improve mixing during sampling, a common pitfall in Energy-based Models. We evaluate KAEM on SVHN and CelebA, demonstrating potential for promising sample quality when correct inductive biases are chosen, while improving interpretability and inference speed against VAEs.
Machine Learning, ICML
@@ 1 Introduction
Latent variable models generate new data from a low-dimensional latent variable passed through a top-down neural generator. Two prominent examples are Variational Autoencoders (VAEs) (Kingma and Welling, 2022) and Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), where the latent vector is typically drawn from a simple, non-informative prior, such as a uniform or isotropic Gaussian distribution. Latent energy-based models (EBMs) (Pang et al., 2020) offer a more flexible alternative by parameterizing and learning a data-dependent prior using a neural network, but their practical adoption remains constrained by the need for Langevin Monte Carlo (LMC) sampling (Brooks et al., 2011; Rossky et al., 1978; Roberts and Stramer, 2002). Given its iterative and gradient-based nature, LMC introduces significant computational overhead and mixes poorly in the multimodal latent distributions common to generative modeling. It also introduces additional experimental complexity since the Unadjusted Langevin Algorithm (ULA) is difficult to tune: step sizes that are too small lead to poor mixing and convergence, while larger step sizes can accumulate discretization bias.
Additionally, despite advances in mechanistic interpretability (Bereska and Gavves, 2024), which seeks to reverse engineer neural networks to understand their internal circuitry, there is no established means of interpreting the trained EBM. This forgoes an opportunity to leverage the latent EBM’s explicit separation between prior and generator for discovery and interpretability, particularly in settings where the underlying structure lies on a low-dimensional manifold that can be identified and represented within the latent space, (e.g, Champion et al. (2019)). Such a feature could enable new ways of embedding inductive bias into generative models, similar to the work of Karniadakis et al. (2021), which integrates scientific biases into neural architectures to reduce data dependence and improve training efficiency, generalization, and interpretability, (key requirements for trustworthy Artificial Intelligence).
A key challenge is therefore to design latent priors that (i) admit efficient and stable inference, (ii) expose interpretable structure, and (iii) can be sculpted with domain knowledge rather than learned purely from optimization dynamics. Current latent EBMs fall short because their priors remain neural parameterizations that are difficult to interpret, slow to sample from, and tightly coupled to stochastic, gradient–based, and iterative sampling methods.
We address these requirements by proposing the Kolmogorov-Arnold Energy Model (KAEM), which revises the latent EBM using a collection of univariate energy functions, based on advances in Kolmogorov-Arnold Networks (KANs) (Liu et al., 2024) and deep spline networks (Bohra et al., 2020). By constraining the prior to univariate relationships, KAEM enables interpretability, alongside fast and exact inference via the inverse transform method for sampling distributions (Devroye, 1986). Additionally, by reinterpreting its inner functions as Markov kernels between probability spaces, we show that KAEM can be viewed as a strict interpretation of the Kolmogorov-Arnold Representation Theorem (KART) (49). Given its finite and deterministic structure, KART can be used as a source of structural bias to reduce the reliance of architecture design on intuition and automated tuning.
@@
KAEM is a uniquely robust probabilistic model, and can enable efficient training with importance sampling (IS) (Goertzel, 1949) on simple datasets. This is a technique for estimating posterior expectations that is otherwise avoided in the literature due to its high variance under severe prior-posterior mismatch and the curse of dimensionality, (Dupuis and Wang, 2004), which can be avoided in our model due to the low dimensionality of the latent space. Moreover, we show that the trained priors are easy to visualize and recover, offering new opportunities for latent discovery and knowledge transfer to models with different latent structures.
Our work additionally introduces novel population-based LMC methods (Swendsen and Wang, 1986; Marinari and Parisi, 1992; Hukushima and Nemoto, 1996) to augment posterior sampling for latent EBMs with a sequence of power posteriors (Friel and Pettitt, 2008) and a training criterion derived from Thermodynamic Integration (Calderhead and Girolami, 2009; Annis et al., 2019). Unlike previous attempts at addressing poor convergence and multimodal exploration in EBMs, which typically rely on diffusion-based models (Cui and Han, 2024; Zhang et al., 2023), the proposed population-based training preserves structure, efficiency, and inference speed. This is because our thermodynamic criterion is solely applied during training and scales in parallel over power posteriors, whereas diffusion-based methods scale sequentially with more denoising steps.
@@
Compilation and differentiation were implemented using Julia’s Reactant (EnzymeAD and Reactant.jl Contributors, 2025) and Enzyme (Moses and Churavy, 2020; Moses et al., 2021, 2022) packages, which optimizes and lowers Julia code to Multi-Level Intermediate Representation (MLIR) (Lattner et al., 2020), thus offering best-in-class performance. However, Gaussian Radial Basis Functions (RBF KANs) (Li, 2024), were adopted instead of the cubic B-spline bases used by Liu et al. (2024) for our interpretability experiments, as RBFs are Reactant-compatible and more efficient for GPU implementation. They have also been proven to provide an optimal solution to the regularized approximation/interpolation problem by Poggio and Girosi (1990). For our VAE comparison experiments, wavelet KANs (Bozorgasl and Chen, 2024) were adopted, since they are better suited to LMC methods as they are not constrained by finite and bounded domains, (unlike RBFs without the grid updating scheme proposed by Liu et al. (2024)).
@@
We evaluate KAEM across several image domains at reduced experimental complexity to demonstrate KAEM’s potential in terms of sample quality and efficiency, as well as the benefit of our proposed training methods. We present KAEM as an initial step to inspire further adaptations of KART in machine learning. This work aims to lay the groundwork for more exploration amongst the research community, with the far-reaching hope of eventually advancing towards a broader conceptualization: “The Kolmogorov-Arnold Representation Theorem Is All You Need.“
[…]
Figure 1: KAEM overview. Generation: Inverse transform sampling (ITS) through a mixture prior enables exact latent sampling without Markov chain Monte Carlo (MCMC). Training: Posterior inference via importance sampling (fast), Langevin dynamics (fast and exploratory), or Thermodynamic Integration (handles multiple modes) provides samples for parameter gradient updates.
2 Preliminaries
The Kolmogorov-Arnold Representation theorem (KART).
@@ 2 Preliminaries
Presented by 49, this forms the basis of KAEM. A structured proof of the theorem has been provided by Dzhenzher and Skopenkov (2022).
Theorem 2.1.
@@
Latent variable generative models.
Let $\bm{x}\in\mathcal{X}$ denote observed data vectors and $\bm{z}\in\mathcal{Z}$ latent vectors. A top-down generative model specifies
$p_{\Phi,f}(\bm{x},\bm{z})\;=\;p_{\Phi}(\bm{x}\mid\bm{z})\,p_{f}(\bm{z}),$ (2)
where $p_{\Phi}(\bm{x}\mid\bm{z})$ is a neural generator (decoder) and $p_{f}(\bm{z})$ is a latent prior parameterized by $f$. The log-marginal likelihood is the learning objective for Maximum Likelihood Estimation (MLE):
$\log p_{\Phi,f}(\bm{x})$ $\;=\;\int\log p_{\Phi,f}(\bm{x},\bm{z})\,p_{\Phi,f}(\bm{z}\mid\bm{x})\,d\bm{z}$
$\;=\;\mathbb{E}_{p_{\Phi,f}(\bm{z}\mid\bm{x})}\left[\log p_{\Phi,f}(\bm{x},\bm{z})\right]$ (3)
Under Appendix. E.0.3, we prove that the resulting learning gradient exhibits a lack of dependence on sampling:
$\nabla_{f,\Phi}\left[\log p_{f,\Phi}(\bm{x})\right]\;=\;$ $\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x})}\Bigg[\nabla_{f}\left[\log p_{f}\left(\bm{z}\right)\right]\Bigg]$
$\;+\;$ $\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x})}\Bigg[\nabla_{\Phi}\left[\log p_{\Phi}\left(\bm{x}\mid\bm{z}\right)\right]\Bigg].$ (4)
Monte Carlo estimator
To approximate arbritrary integrals such as the expectation in Eq. 4, one can draw $N$ samples from the target distribution and take an unbiased Monte Carlo estimator. For some arbritrary function, $\rho$, (e.g., the log-likelihood), the expectation with respect to an arbritrary density, $p$, can be estimated as:
$\mathbb{E}_{p(z)}[\rho(z)]\;=\;\int_{\mathcal{Z}}\rho(z)\;p(z)\;dz\;$ $\approx\;\frac{1}{N}\sum_{s=1}^{N}\rho(z^{(s)}),$
$\text{where}\quad\{z^{(s)}\}_{1:N}$ $\;\sim\;p(z).$ (5)
Top-down generator
The latent-to-data generative process is realized using a neural network $G_{\Phi}:\mathcal{Z}\to\mathcal{X}$, (parameterized by $\Phi$), which maps a latent sample $z$ to a reconstruction (mean image or data sample)
$\tilde{\bm{x}}\;=\;\mu_{\Phi}(\bm{z})\;=\;G_{\Phi}(\bm{z})+\bm{\varepsilon},\qquad\bm{\varepsilon}\sim\mathcal{N}(\mathbf{0},\sigma_{\text{noise}}\mathbf{I}),$ (6)
In this work we model the conditional likelihood with an isotropic Gaussian observation model, equivalent to a pixel-wise $\ell_{2}$ reconstruction objective:
$p_{\Phi}(\bm{x}\mid\bm{z})\;=\;\mathcal{N}\!\big(\bm{x};\,\mu_{\Phi}(\bm{z}),\,\sigma_{x}^{2}\mathbf{I}\big),$ (7)
where $\sigma_{x}^{2}$ is a specified variance, $\sigma_{\text{noise}}^{2}$ captures observation noise, and $\mathbf{I}$ is the identity matrix over pixels and channels. For an image $x\in\mathbb{R}^{H\times W\times C}$, the corresponding log-likelihood is
$\log p_{\Phi}(\bm{x}\mid\bm{z})\;\propto\;-\frac{1}{2\sigma_{x}^{2}}\,\|\bm{x}-\mu_{\Phi}(\bm{z})\|_{2}^{2}$ (8)
Up to an additive constant, maximizing $\log p_{\Phi}(\bm{x}\mid\bm{z})$ is equivalent to minimizing the pixel-wise squared error. The gradient needed for the log-likelihood component of Eq. 4 can then be found by autodifferentiation, which we used Enzyme (Moses and Churavy, 2020) for.
Latent energy-based prior.
Latent EBMs (Pang et al., 2020) define the prior by an energy $E_{f}$ and base reference prior, e.g., $\pi_{0}=\mathcal{N}(0,1)$:
$p_{f}(\bm{z})\propto\exp(E_{f}(z))\cdot\pi_{0}(z).$ (9)
Training requires sampling from the posterior
$p_{\Phi,f}(\bm{z}\mid\bm{x})\propto p_{\Phi}(\bm{x}\mid\bm{z})\,p_{f}(\bm{z}),$ (10)
which is often multimodal and difficult to sample from.
Unadjusted Langevin Algorithm (ULA).
Given a target $\pi(z)\propto\exp(\log\pi(z))$, ULA (Brooks et al., 2011) iterates
$z^{i+1}=z^{i}+\frac{\eta}{2}\nabla_{z}\log\pi(z^{i})+\sqrt{\eta}\,\xi^{i},\quad\xi^{i}\sim\mathcal{N}(0,I),$ (11)
with step size $\eta>0$. In latent EBMs, $\pi$ is either the prior (9) or posterior (10). ULA is best suited to sampling from unimodal distributions. Following Pang et al. (2020) we adopt a small iteration count, (40 iterations), following the guidance of Nijkamp et al. (2019b, a) regarding short-run Markov chain Monte Carlo, and to limit the resource expenditure of our experiments.
Power posteriors.
To improve mixing, KAEM samples from power posteriors (Friel and Pettitt, 2008):
$p_{\Phi,f,t}(\bm{z}\mid\bm{x})\propto p_{\Phi}(\bm{x}\mid\bm{z})^{t}\,p_{f}(\bm{z}),\qquad t\in[0,1],$ (12)
where $t=0$ recovers the prior and $t=1$ recovers the posterior. A population of chains at $\{t_{k}\}_{k=0}^{K}$ with swap moves (parallel tempering) improves exploration. Thermodynamic Integration connects expectations under $p_{t_{k}}$ to log-normalizer differences; details are deferred to Sec. 3.2.
3 Method
3.1 Architecture
KAEM prior.